OpenStudy (anonymous):

Solve for x: (x – 4)2 = 2 x = 4 ± 2 sqrt2 x = 4 ± sqrt13 x = 4 ± sqrt2 x = 4 ± 2 sqrt13

OpenStudy (ash2326):

We have \[(x-4)^2=2\] take root both the sides, we get \[x-4=\pm \sqrt 2\] Now add 4 to both the sides @kenbaby

OpenStudy (jiteshmeghwal9):

use identity\[(a-b)^2=a^2+b^2-2ab\]so,\[x^2+16-8x=2\]\[x^2+16-2-8x=0\]\[x^2+14-8x=0\]